S + 5 3. (4pts) Find L-1 $2 + 6s + 18
Answer #1

solution

L^{-1}\left\{\frac{s+5}{s^2+6s+18}\right\}\\ \mathrm{Expand\:}\frac{s+5}{s^2+6s+18}:\quad \frac{s+3}{\left(s+3\right)^2+9}+2\cdot \frac{1}{\left(s+3\right)^2+9}\\=L^{-1}\left\{\frac{s+3}{\left(s+3\right)^2+9}+2\cdot \frac{1}{\left(s+3\right)^2+9}\right\}

by using linearity property of inverse laplace transform, we get

=L^{-1}\left\{\frac{s+3}{\left(s+3\right)^2+9}\right\}+2L^{-1}\left\{\frac{1}{\left(s+3\right)^2+9}\right\}\\L^{-1}\left\{\frac{s+3}{\left(s+3\right)^2+9}\right\}:\quad e^{-3t}\cos \left(3t\right) (using formula)

L-1{s+a/(s+a)2+b2}=e-at.cos(bt)

L^{-1}\left\{\frac{1}{\left(s+3\right)^2+9}\right\}:\quad \frac{1}{3}e^{-3t}\sin \left(3t\right)\\=e^{-3t}\cos \left(3t\right)+2e^{-3t}\frac{1}{3}\sin \left(3t\right)\\=e^{-3t}\cos \left(3t\right)+\frac{2}{3}e^{-3t}\sin \left(3t\right) .. answer

=e-3t(cos(3t)+2/3.sin(3t))